OpenStudy (anonymous):

What is the average acceleration of a jet ski that goes from rest to 10 m/s west in 4 s? Take the direction west to be positive, and east to be negative. A. 2.5 m/s2 B. -2.5 m/s2 C. 0.4 m/s2 D. -0.4 m/s2

OpenStudy (danjs):

THe average acceleration is the slope of a secant line connecting those 2 given points

OpenStudy (danjs):

(v(4) - v(0)) / ( 4 - 0 )

OpenStudy (anonymous):

uh?

OpenStudy (danjs):

v(4) = 10 m/s v(0) = 0 m/s rest

OpenStudy (danjs):

\[<a> = \frac{ v(4) - v(0) }{ 4-0 } = \frac{ 10 - 0}{ 4 - 0 } ~m/s^2\]

OpenStudy (anonymous):

Uh sorry, i just dont understand, just too confused. A = ? V = ? M/S = ? (Meters per second?) Anyway im just really running out of time so if someone could just help me asap it would really save me!